Use Cauchy’s integral formula to evaluate sin(z) /((2z + 1)*3e*z) dz


1
Expert's answer
2021-11-21T11:55:21-0500

Cauchy’s integral formula:


"f(z_0)=\frac{1}{2\pi i}\int_C \frac{f(z)}{z-z_0}dz"


then:


"\frac{sinz}{ 3ez(2z + 1)}=\frac{sinz}{ 3e}\cdot\frac{1}{ z(2z + 1)}"


"\frac{1}{ z(2z + 1)}=\frac{1}{ z}-\frac{2}{ 2z + 1}"


"\int_C \frac{sinz}{ 3ez(2z + 1)}dz=\int_C \frac{sinz}{ 3ez}dz-\int_C \frac{2sinz}{ 3e(2z + 1)}dz"


for "\int_C \frac{sinz}{ 3ez}dz" :


"f(z)=\frac{sinz}{ 3e},z_0=0"


"f(0)=0\implies \int_C \frac{sinz}{ 3ez}dz=0"


for "\int_C \frac{2sinz}{ 3e(2z + 1)}dz" :


"f(z)=\frac{sinz}{ 3e},z_0=-1/2"


"f(-1\/2)=-0.059"


so,


"\int_C \frac{sinz}{ 3ez(2z + 1)}dz=0.059\cdot2\pi i =0.369i"